# Rise and fall of a multicomponent droplet in a surrounrdfing fluid: simulation study of a bumpy path

### v1

@@ III Simulation setup
The uniform mesh size is set to 2×$\times 10^{-4}$ m. As a result there are 4 cells within the interface so that the interface structure can be captured and that grid anisotropy can be neglected. Accordingly, the upscaling parameter $\lambda$ is introduced in the phase field model in order to treat a larger interface thickness than the characteristic length scale of the physical interface Cardon et al. (2016b); Rasolofomanana et al. (2022a). With such a thermodynamic landscape and the related equilibrium compositions, the parameters $\kappa^{bin}$ and $\lambda$ can be defined from the surface tension $\sigma$ and interface thickness $\varepsilon$ as:
@@
$\phi_{i}(x,y,t=0)=\dfrac{\phi_{i}^{ini,dis}}{2}\left(1-\tanh\left(\dfrac{\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}-R}{\varepsilon}\right)\right)$ (11)
@@
Table 2: Summary of the compositions of elements A and B with the density of each phase at initial and equilibrium states.
phase initial state equilibrium state
dispersed $\phi_{A}^{ini,dis}=0.15$ $\phi_{A}^{eq,dis}=0$
phase $\phi_{B}^{ini,dis}=0.85$ $\phi_{B}^{eq,dis}=1$
(droplet) $\rho_{dis}^{ini}=980$kg.m-3 $\rho_{dis}^{eq}=1050$kg.m-3
continuous $\phi_{A}^{ini,cont}=0$ $\phi_{A}^{eq,cont}=\frac{\phi_{A}^{ini,dis}V_{dis}^{ini}}{V_{tot}+(\phi_{A}^{ini,dis}-1)V_{dis}^{ini}}$
phase $\phi_{B}^{ini,cont}=0$ $\phi_{B}^{eq,cont}=0$
$\rho_{cont}^{ini}=999.5$kg.m-3 $\rho_{cont}^{eq}=\rho_{wat}\left(1+\beta_{A}\phi_{A}^{eq,cont}\right)$
Before turning to the result, we find necessary to discuss how some quantities related to the droplet are computed. Indeed, since there is practically no mass transfer of immiscible element to the continuous phase, we are particularly interested in the amount of immiscible element in presence. Thus, we can define the position of the droplet from the position of the center of mass and its velocity by:
@@
where $\phi_{B}$ is the molar fraction of the immiscible component, $\bm{x}(x,y)$ and $\bm{u}$ are the spatial coordinates and velocities of a point.
IV Results and discussion
@@
Figure 2: Position of the droplet as a function of time for different values of the mobility.
We now present numerical results obtained with this model and discuss them. First, we present the results from a qualitative point of view. Second, we discuss the interplay between diffusive mass transport and advection and show that it has significant effects on the droplet dynamic that can not be neglected. This emphasises the importance of fully coupled numerical simulations as the one presented here.
IV.1 General behaviour
@@
As expected, calculations show that $x_{G}$ remains constant throughout the simulation. That implies that the droplet follows a rectilinear trajectory as shown in fig 2 where the position of the droplet along the vertical axis $y$ is plotted as a function of time for different values of $M$. Its trajectory includes an ascending phase followed by a descending phase. In the ascending phase, the droplet being lighter than the surrounding water first starts to rise. In the same time, the miscible component of the droplet diffuses out in water, resulting in an increase of the droplet density. After a while, the droplet reaches a maximum height, becomes heavier than the continuous phase and ends up falling down into the water. This behaviour is in good qualitative agreement with experimental results presented in Rao et al. (2015) and is robust to changes in parameters such as the mobility. However, one can see that changes in the mobility dramatically affect the amplitude and duration of its motion. For instance, for $M=10^{-8}J^{-1}m^{5}s^{-1}$, the amplitude of the droplet motion is $\sim R/2$ (R is the radius of the droplet) while for $M=5\times 10^{-10}J^{-1}m^{5}s^{-1}$, it corresponds to $\sim 20R$. This is consistent with the fact that the characteristic diffusion time $\tau$ obtained by scaling the Cahn-Hilliard equations is proportional to the inverse of the mobility i.e. $\tau\sim\dfrac{1}{M}$. For lower mobilities, the diffusion is slower. Hence, the droplet rises high enough. On the contrary, when mobility $M$ is greater, the diffusion becomes faster and the droplet is unable to reach the same height.
Figure 3: Typical evolution of the droplet velocity as a function of time where the density maps are also shown at different times. The mobility is $5.10^{-9}$.
In order to analyse the droplet motion, in the following, we will use the droplet velocity evolution curve that allows a better analysis. Therefore, we first give a qualitative description of the evolution of the droplet velocity.
A typical curve of the velocity along the vertical axis together with maps of the density at given times is presented in Fig. 3 for $M=5\times 10^{-9}J^{-1}m^{5}s^{-1}$. One can see that after an acceleration phase that is absent from experimental curves in Rao et al. (2015), the velocity of the droplet reaches a maximum and starts to decrease linearly (with a few bumps) until it reaches a limiting velocity. When looking at the density map, with the exception of the initial map, the boundary between the droplet and the continuous phase appears clearly (in supplemental material we present the corresponding maps of all fields for the cases presented here at different times). It can be seen at the two intermediate steps that there is a light fluid layer (A rich without any B) around the droplet that is attached to the droplet when the droplet velocity is maximum and that has partially detached from the droplet when its vertical velocity cancels.
It is worth mentioning that when the droplet velocity is at its maximal value, the droplet density is already larger than the density of water which is unexpected. This can be explained by two facts. First, inertia introduces a delay (this will be discussed in the next section), second, if one consider the volume occupied by the droplet and the surrounding diffusive layer, its average density is the average density of the initial droplet and of the surrounding pure water. Since the initial droplet was lighter than water, this average density is smaller than the density of water and this virtual volume of fluid is pushed upward by buoyant forces.
Thereafter, the droplet velocity decreases, but the droplet is still rising. One can see that the surrounding layer of light fluid has detached from the droplet due to buoyant forces. At this stage, the droplet is significantly denser than water and its velocity cancels. The droplet falls in the fluid leaving behind a plume of A rich water until the droplet is composed of pure B and reaches a limiting velocity.
@@
These observations already illustrate the interplay between diffusion and flow. Indeed, one can see that the droplet density is asymmetric and that it is denser in its lower part. This is related to the fact that, as can be seen on the map, the diffusive layer is thinner at the bottom of the droplet than at its top. This leads to a more efficient diffusive transport of A out of the droplet, a faster decrease in A at the bottom of the droplet, which translates into a heavier fluid. This configuration is stable with respect to buoyant forces, however, its origin comes from the interplay of diffusion and flow and not some sedimentation.
In the following section, the effect of the mobility on the droplet motion is discussed.
IV.2 Effect of the mobility
@@
In the case of a quiescent droplet, the evolution of its composition is governed by Eq. 1 with the mobility $\bar{\bar{\mathcal{M}}}$ given by Eq. 3 and a velocity $\bm{u}=0$. It can be rewritten as
$\dfrac{1}{M}\dfrac{\partial\phi_{i}}{\partial t}=\nabla\cdot(\nabla\tilde{\mu}_{i})$ (14)
and $1/M$ is proportional to the characteristic time of evolution of the composition of the droplet. This implies that for a quiescent droplet with varying values of $M$, the exact same behaviour is expected after a rescaling of time by $1/M$. Since the droplet motion affects the diffusion, we do not expect this scaling to hold strictly speaking. However, for high values of the mobility (for low values of the buoyant forces, which is not studied here, this scaling is also valid), it should be the main effect of mobility. Since inertial effects cannot be neglected, we do not expect that the curves of velocity as a function of rescaled time superimpose for high values of mobility. This is indeed the case, as can be seen in Fig. 4 where the droplet velocity is plotted as a function of the time rescaled by $M$.
Figure 5: Molar fraction field of the element A at $t\times M=1.5\times 10^{-9}J^{-1}m^{5}s^{-1}$ for different mobilities $M$ using either a linear scale (first column) or a logarithmic scale (rightmost column). The bubble defined as an isoline of B element is drawn and the scale between the columns is not the same. The white bar corresponds to $\approx 3.mm$ Only a portion of the domain is shown in both cases.
Before discussing the results and the peculiar shape of the curves, we discuss a simple phenomenological model that aims at putting under light the simple effects of both inertia and diffusive transport. To this purpose, we consider a simple model of the droplet motion:
$m\dfrac{du}{dt}=f(Mt)-\lambda u$ (15)
where $m$ is an effective mass, $f(Mt)$ is a function that describes the effect of buoyant forces and that evolves due do diffusive process and $\lambda$ is an effective friction. We consider the simple case where $f(t)=a(1-t/t_{0})$. Taking into account the initial condition $u(0)=0$, this simple model can be solved analytically and has for solution
$u(t)=\dfrac{a}{\lambda}\left[\left(1+\dfrac{mM}{\lambda\xi_{0}}\right)\left(1-e^{-\lambda t/m}\right)-\dfrac{Mt}{\xi_{0}}\right]$ (16)
Using $a=0.035$m/s, $\xi_{0}=2.5\times 10^{-9}$-1m5, $m=0.04$kg and $\lambda=0.08$kg/s, one can get the curves presented in Fig. 4. Parameters were chosen so that the time at which the velocity cancels for the highest mobility is approximatively the same as in numerics and the peak velocity for the lowest mobility is also in agreement with numerics. It must be noted that with such values of the parameters, the buoyant force changes sign at $t\times M=2.5\times 10^{-9}J^{-1}m^{5}$ in the model.
Figure 6: Velocity vectors in the frame of the droplet in the density field for $M=2.5\times 10^{-9}J^{-1}m^{5}s^{-1}$ at different times that correspond to the bump in the velocity field. Light regions are represented in yellow while dense regions in red-purple. Pure water density is orange. Only a portion of the domain si shown.
In the curve obtained with the model for the two lowest mobilities, the velocity changes sign at $t\times M\approx 2.5\times 10^{-9}$, indicating that in these two cases, once the initial acceleration phase is over, the driving force time scale is larger compared to the inertial time scale, so that the later can be neglected. On the opposite, for $M=10^{-8}J^{-1}m^{5}s^{-1}$, the maximum of the velocity corresponds to $t\times M\approx 2.5\times 10^{-9}$.
When comparing with reduced model results, one can see that the general shape of the velocity curve is in good agreement with numerics and the general qualitative behaviour is recovered. However, in the case of low mobility values, a few differences can be seen.
For $M=5\times 10^{-10}J^{-1}m^{5}s^{-1}$, one can see that the main difference is that the time at which the velocity cancels is slightly smaller in the numerics than in the reduced model. This difference is of the order of approximately 20% and is related to the advective transport of the A component leading to a faster evolution. In addition, one can see a small bump in the numerical velocity curve close to its maximum. This aforementioned velocity bump is also present (and much more visible since it translates into a well marked velocity plateau) in numerical results for $M=1,2.5\mbox{ and }5\times 10^{-9}J^{-1}m^{5}s^{-1}$. We give a rationale for these two effects in the following.
First, the fastest evolution of the droplet velocity for small values of $M$ can be attributed to the fact that due to the advection, the droplet is moving in fresh water. As a result, the diffusion layer is small and the diffusive transport is fast, leading to a faster transport of $A$ outside of the droplet. This can be seen in Fig. 5 where the composition map of A is represented for different values of $M$ at rescaled time $M\times t=10^{-9}J^{-1}m^{5}$. Regardless of $M$ ($M\in[5\times 10^{-10},10^{-8}]J^{-1}m^{5}s^{-1}$), the droplet is in its ascending phase at this value of $M\times t$. However, the distribution of the miscible element in the droplet and its diffusion in water are different depending on the value of $M$. For larger values of $M$ (i.e. $M=\{5\times 10^{-9},10^{-8}\}J^{-1}m^{5}s^{-1}$), the diffusion dominates and the droplet behaves in almost the same way for these two values. On the other hand, its behaviour is significantly different for the lowest mobilities $M\leq 2.5\times 10^{-9}J^{-1}m^{5}s^{-1}$ where diffusion is slower. The diffusion field is then modified by the flow. This can be seen in Fig. 5 where the volume fraction of element A is represented using pseudo colors at $t\times M=1.5\times 10^{-9}J^{-1}m^{5}$ for different mobilities using both linear and logarithmic scales. One can clearly see that the extent of the diffusion layer is asymmetrical in the low mobility case (b1 and c1) while it has almost cylindrical symmetry in the high mobility case (a1). In addition, using the linear scale, one can see that, the diffusion layer is thinner both ahead and behind the droplet leading to faster diffusive transfer for low mobilities.
Second, we discuss the presence of a plateau in the velocity curve for small values of the mobility. Since it is well marked for $M=2.5\times 10^{-9}J^{-1}m^{5}s^{-1}$ we will focus on this case. First one should notice it has significant consequences: the rescaled time at which the droplet starts to sink (cancellation of the velocity) is significantly larger than in the high mobility case while, as explained above, the advection alone would lead to an earlier descent. The plot of the velocity field in the frame of the droplet (that is the velocity field in the laboratory frame minus the average velocity of the droplet) together with the fluid density in Fig. 6 allows to better understand this phenomenon. All plots correspond to the ascending phase. The first one is well before the bump and one can see a droplet moving in a flow that is very similar to a Stokes flow around an obstacle (the droplet is moving from left to right). The second one corresponds to a time close to the onset of the plateau and one can see that next to the droplet, there is a boundary layer close to the droplet that is moving faster than it in the same direction. It corresponds to the region of the surrounding fluid that is rich in A element and is lighter and it induces a viscous force in the same direction as the droplet motion. For later times, this effect is stronger, however, one can see that the light region is moving slowly toward the front of the droplet and eventually detaches from the droplet, leading to the end of the plateau. This effect is similar to the drafting technique in a peloton Belloli et al. (2016); Marro et al. (2023).
The further evolution of the droplet is simple, it keeps losing A component until it reaches a terminal velocity that corresponds to the Stokes velocity of a droplet of pure B in a water bath.
V Conclusion

### v2

@@ IV Simulation setup
The upscaling parameter $\lambda$ in the phase field model allows to consider a larger interface thickness than the characteristic length scale of the physical interface Cardon et al. (2016); Rasolofomanana et al. (2022a). With such a thermodynamic landscape and the related equilibrium compositions, the parameters $\kappa^{bin}$ and $\lambda$ can be defined from the surface tension $\sigma$ and interface thickness $\varepsilon$ as:
@@
The numerical results are expected to converge toward the actual solution when $\varepsilon\to 0$ provided that the mobility $M$ is properly chosen, as will be discussed later.
@@
$\phi_{i}(x,y,t=0)=\dfrac{\phi_{i}^{ini}}{2}\left(1-\tanh\left(\dfrac{\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}-R}{\varepsilon}\right)\right)$ (11)
@@
The main quantities of interest in these simulations are the related to the droplet motion. Accordingly, since no mass transfer of the immiscible component is expected we define the position of the droplet and its velocity by:
@@
where $\phi_{B}$ is the molar fraction of the immiscible component, $\bm{x}(x,y)$ and $\bm{u}$ are the spatial coordinates and velocities of a point. In the next section we discuss in more details the choice of parameters and the convergence of the model.
V Choice of parameters and convergence
@@
Figure 2: position of the droplet center of mass along the $y$ axis as a function of $t/\tau_{d}=8.33$ (see eq.15) for different values of the interface thickness $\varepsilon$.
First we have checked the sensibility of our simulations to grid spacing and have found that using $dx=\varepsilon/4$ lead to converged results with respect to grid spacing: doubling the resolution did not change the results and the curves of the droplet position as a function of time for both values of $dx$ superimposed. We have used this resolution throughout our simulations.
Then we have checked the convergence of our simulations with respect to the interface thickness. To this purpose, we need to discuss the choice of the mobility $M$. Here the droplet motion evolution is due to diffusion coupled to Navier-Stokes equation. Using the assumption that the interface is at equilibrium, the process that governs the motion is the diffusive mass transport far from the interface. In this region (far away from interface), the Laplacian term of the chemical potential can be neglected and the transport equation at a given point , using a second order Taylor expansion of $G$ around the concentration field at this point (with a diagonal $\mathcal{M}$) writes:
$\frac{\partial\phi_{i}}{\partial t}\approx M\Delta\lambda\frac{dG}{d\phi_{i}}=(M\lambda)\left(\frac{d^{2}G}{d\phi_{i}^{2}}\right)_{\phi_{i}}\Delta\phi_{i}=D_{i}\Delta\phi_{i}$ (14)
As a result, to keep the bulk diffusion kinetic independent of the interface thickness $M\lambda\propto M/\varepsilon$ is kept constant when varying the interface thickness, keeping the surface tension $\sigma$ constant. With this choice of parameters we have checked the convergence of the simulations with respect to the interface thickness. To this purpose we have performed numerical simulations with the interface thickness ranging from 0.0001 to 0.0008 and, as can be seen in fig. 2 where the position of the droplet for $D_{A}=1.08.10^{-6}m^{2}s^{-1}$ is plotted as a function of time. This value of $D_{A}$ corresponds to a diffusion time $\tau_{d}=8.33s$ defined as
$\tau_{d}=\frac{l^{2}}{D_{A}}$ (15)
[…]
It must be noted that the curve obtained for $\varepsilon=0.004$ is very similar (qualitatively and quantitively) to the curve obtained for $\varepsilon=0.002$ and $\tau_{d}=11.11$ (not shown). Hence the lack of convergence could be related to an increased interfacial resistance when the interface thickness is increased that is sufficiently high to slow down the diffusive transport. A similar behaviour is observed for heat transport in the case of rough interfaces Merabia and Termentzidis (2014).
[…]
$M_{\varepsilon=0.0008}$ $M_{\varepsilon=0.0002}$ $D_{A}$ $\tau_{d}$
100.e-10 25.e-10 1.08e-5 0.833
50.e-10 12.5e-10 5.4e-6 1.666
20.e-10 5.e-10 2.16e-6 4.166
10.e-10 2.5e-10 1.08e-6 8.333
5.e-10 1.25e-10 5.4e-7 16.66
2.e-10 0.5e-10 2.16e-7 41.66
1.e-10 0.25e-10 1.08e-7 83.33
Table 2: Table of the mobilities for two values of the interface thickness and the corresponding values of the diffusion constant $D_{A}$ and the corresponding diffusive characteristic time $\tau_{d}$. Units for diffusion constants are $m^{2}/s$, for $\tau_{d}$ they are seconds and mobilities are expressed in.
VI Results and discussion
Figure 3: Position (left) and speed (right) of the droplet as a function of time for $\varepsilon=0.0002$ and $\tau_{d}=8.333$.
We now present numerical results obtained with this model and discuss them.
VI.1 General behaviour
@@
As expected, the position of the droplet center along the horizontal axis remains constant throughout the simulation. In fig 3 the position of the droplet along the vertical axis $y$ is plotted as a function of time for $\tau_{d}=8.33s$. Its trajectory includes an ascending phase followed by a descending phase. In the ascending phase, the droplet being lighter than the surrounding water first starts to rise. In the same time, the miscible component of the droplet diffuses out in water, resulting in an increase of the droplet density. Eventually, the droplet reaches a maximum height, becomes heavier than the continuous phase and ends up falling down into the water. This behaviour is in good qualitative agreement with experimental results presented in Rao et al. (2015). Due to the nature of our simulations and the value of parameters used here, quantitative agreement cannot be reached and this is out of the scope of this paper. Indeed, our simulations have been performed in 2D geometry while experiments are 3D. In addition the initial bubble composition used here differs significantly from the one in Rao et al. (2015) and its initial mass density is 980kg.m-3, which is heavier than in experiments. Some more information about the comparison between such Cahn-Hilliard based simulations (carried out with TrioCFD) and the experimental results can be found in Rasolofomanana (2023).
Figure 4: Colormaps of the fluid density (top row) and of the A component concentration (bottom row) at different rescaled times $t/\tau_{d}$(the values label the density maps). The droplet (the B rich region) is shown by the green circle.. Parameters are the same as in figs. 3. The structure of the flow is illustrated by the flow field with respect on the top row in the frame. The represented field is $\mathbf{u}-v_{y}\mathbf{e_{y}}$ where $v_{y}$ is the vertical velocity of the droplet.
We present a typical curve of the velocity along the vertical axis in fig. 3 for $\tau_{d}=8.333$. After an acceleration phase, that is absent from experimental curves in Rao et al. (2015), the velocity of the droplet reaches a maximum and starts to decrease. However contrarily to experimental results the decrease of the velocity is not as regular as it would be expected from a naïve approach. Indeed, the evolution of the droplet velocity is clearly non monotonous and presents a bump before a fast acceleration of the droplets leads to its fall (it eventually reaches a terminal velocity). In order to better understand this unexpected bump we present concentration maps of $A$ together with the flow field and fluid density maps at different times of the evolution of the droplet in fig. 4. The colormaps are taken at the beginning of the motion, when the droplet velocity reaches its maximal value, after it has decreased and reaches a first minimum and close to the moment of the bump in the droplet velocity. Hence the discussion here is focused on the upward motion of the droplet.
The first observation is that the droplet presents in heterogeneous concentration in $A$ with a minimum close to its bottom. In addition, one can see that at the interface there is a maximum of A concentration that is similar to the bumps that can be seen in some concentration profiles at equilibrium Rasolofomanana et al. (2022a). Before going further we want to discuss the issue of these bumps. Since the thermodynamic landscape has been chosen to avoid such bumps they do not share the same cause as the one observed in Rasolofomanana et al. (2022a). An analysis using the chemical potential of $A$ shows that the concentration bump do not correspond to any chemical potential bump and that when going through the interface, the gradient of chemical potential is constant which corresponds to a constant mass flux when crossing the interface as expected. Hence these bumps are simply related to the fact that as mentioned in Plapp (2011) when going through an interface the quantity that is continuous is not the concentration of a solute but its chemical potential. It must be noted that outside of the interface, the concentration of $A$ and its chemical potential are proportional to each other.
The higher concentration of A at the top of the droplet is in correlation with the fact that the $A$ diffusion layer is larger at the top of the droplet than at its bottom. This asymmetry of the diffusion layer is in contradiction with what is expected when considering the forced movement of an object that is releasing some chemical in the surrounding fluid: in this case the upper part of the droplet is advancing in a fresh fluid and the diffusion layer ahead of the droplet should be thinner than behind it. Here the opposite is observed. This is due to the fact that the flow is affecting the diffusive layer. Indeed, since $A$ is lighter than water (and than the droplet), there is an upward motion of the diffusion layer with respect to the drop that leads to the formation of the thicker layer of $A$ above. This illustrated by the flow pattern where one can see that close to the droplet the flow is moving upward with respect to the droplet due to buoyant forces.
@@
These observations already illustrate the effect of diffusion on flow. On the density map, one can also see effects of the flow on diffusion. Indeed, the droplet density is asymmetric and that it is denser in its lower part. This is related to the fact that, as seen on the map, the diffusive layer is thinner at the bottom of the droplet than at its top. This leads to a more efficient diffusive transport of A out of the droplet, a faster decrease in A at the bottom of the droplet, which translates into a heavier fluid. This configuration is stable with respect to buoyant forces, however, its origin is the interplay between diffusion and flow and not some sedimentation. It must be noted that the asymmetry of the diffusion layer is the opposite of what is expected if there is no effect of the concentrations on flow. One would expect a larger concentration of $A$ behind the droplet.
VI.2 Effect of the mobility
In order to illustrate the effects of the interplay between flow and diffusion we find it interesting to first consider a system where such effects are absent. To this purpose we consider a simple model that describes the inertial motion of an object that is gaining mass along time and that is subject to a viscous drag force. The equation of motion of this object writes:
$\dfrac{du}{dt}=f(t)-\frac{u}{t_{r}}$ (16)
where $f(t)$ is a function that describes the effect of buoyant forces and that evolves due to diffusive process and $t_{r}$ is the characteristic time associated with the effective viscous drag. We consider the simple case where the driving force for the motion $f(t)=\frac{v_{0}}{t_{r}}(1-t/t_{0})$, with $v_{0}$ a characteristic velocity. With this expression, $t_{0}$ is a characteristic time that is proportional to $\tau_{d}$. Taking into account the initial condition $u(0)=0$, this simple model can be solved analytically and has for solution :
$u(t)=v_{0}\left[\left(1+\dfrac{t_{r}}{t_{0}}\right)\left(1-e^{-t/t_{r}}\right)-\dfrac{t}{t_{0}}\right]$ (17)
We have chosen to determine the model parameters with a fit of the early stage (i.e. the acceleration part and the onset of the slowdown) of simulations using an intermediate value of $\tau_{d}=8.333s$. A good fit is reached for the following parameter values: $t_{0}=0.806$, $v_{0}=0.056$ and $t_{r}=0.709$. Using these values and considering that $t_{0}$ is proportional to the diffusion time, we plot the curves of the velocity as a function of time rescaled by $\tau_{d}$ for different values of $\tau_{d}$ on the left panel of fig. 5. On the bottom panel the actual numerical results are plotted. There is a significant discrepancy between the results of the reduced model and the actual numerical results. It is simply quantitative for high and low values of the characteristic time for diffusion $\tau_{d}$. But for intermediate values the simple model does not capture the bump in the velocity curve that is seen and has already been discussed in the previous section. This bump has significant effects since it corresponds to a significant increase in the maximum height reached by the droplet. This can be seen in fig. 6 where the maximum height is plotted as a function of the diffusion time.
@@ VI.2 Effect of the mobility
For low mobilities (high $\tau_{d}$), the maximum height scales linearly with $\tau_{d}$, which is expected from a simple scaling argument since $\tau_{r}<<\tau_{d}$. For intermediate mobilities, the maximum height is significantly higher than this linear scaling (up to 2.5 times higher). While for low values of $\tau_{d}$ (smaller than $t_{r}$) the maximum height scales like $\tau_{d}^{2}$ due to inertial effects: the times it takes for the bubble to accelerate is larger than $\tau_{d}$.
Figure 6: left Position of the maximum $y$ of the trajectory of the droplet as a function of the diffusion time $\tau_{d}$. The line, of equation $y=\alpha\tau_{d}$ with $\alpha=0.002222$ corresponds to a linear fit of the data for the lowest mobilities. right Ratio of the maximum $y$ of the trajectory with the fit for low mobilities.
Finally, as its importance on the motion of the droplet has been made obvious here, we find it useful to discuss the diffusion layer around the droplet for different values of the mobility. To this purpose, we have plotted the diffusion layer for $\tau_{d}=8.33,\ 41.66\mbox{ and }1.66s$ at $t/\tau_{d}=0.06,\ 0.12\mbox{ and }0.18$ in fig. 7. From these maps, one can see first that for slow diffusion, the flow leads to the thinning of the diffusion layer around the droplet (1st line). This in turn leads to higher gradients in the bulk and therefore to higher A flux outward of the droplet. As a result, as can be seen on the last line, for higher mobility the amount of A in the droplet at a given rescaled time is higher. When one consider the diffusion layer, it is also clear that it presents a vertical asymmetry: it is thicker above the droplet (which is moving upward) while in the case of a passive diffusing component it should be thicker below. This illustrate the fact that the lighter A rich fluid around the droplet is moving upward. As a result the diffusive transport of $A$ is weaker in the region above the droplet and there is an higher concentration of $A$ in the top part of the droplet.
This translates into solutal Marangoni forces that affect the motion of the fluid. These forces are due to the $\Sigma\tilde{\mu}_{i}\nabla\phi_{i}$ term in eq. 4 that is the term that leads to the surface tension. This vector field is represented in fig. 8 a together with the concentration map of the miscible component and one can see that it leads to mostly inwardly pointing vectors located at the interface. In order to keep the figure readable the vectors are plotted only one every 8 point in each direction. In fig. 8 b, to represent Marangoni forces these vectors are plotted after a projection on the direction tangential to the interface (the gradient of $\phi_{B}$ ) and a 25 times scaling. One can see that there is a component tangential to the interface that would correspond to a Marangoni force that tends to push the interface upward (which would reinforce the effect of the buoyant layer. However, as can be seen some of the represented vectors show a discontinuity. This is due to the fact that for vanishing gradients of $\phi_{B}$ (in the vicinity of the interface) the projection operator is discontinuous. In addition with this approach the Marangoni forces are also present in the bulk and not located at the interface. Therefore while our analysis indicates that solutal Marangoni forces are present and are contributing to the droplet motion, it is difficult to quantify their magnitude when compared to other sources of motion. In less complicated context, such as the bouncing droplet Li et al. (2019), we believe that our model could provide valuable insight.
Figure 7: colormaps of the A concentration for three different diffusion times ($\tau_{d}=8.33,\ 41.66\mbox{ and }1.66$ ) at three different dimensionless times $t/\tau_{d}=0.06,\ 0.12\mbox{ and }0.18$. The colorscale is the same as in fig. 4.
Figure 8: a Colormaps of the A concentration together with the vector plot of source term for surface tension. b same plot with the source term projected along the tangent to the interface (i.e. normal to $\nabla\phi_{B}$) and magnified by a factor of 25 when compared to a. The parameters are $\tau_{d}=8.333$ and $t/\tau_{d}=0.096$. At this time the droplet is moving upward. The orientation of the gravity field is given by the arrow.
VII Conclusion
We have carried out a 2D simulation of the dynamics of a multicomponent droplet released in a water column using a ternary Cahn-Hilliard Navier-Stokes model. This model describes naturally coupled mass transfer and hydrodynamic phenomena. To illustrate both the capacity of the model to solve such problem and the interest of fully coupled simulations we have considered the case of a multicomponent droplet rising and sinking in water. We have first studied the model convergence with respect to the interface thickness and shown that for interface thickness less than $1/15$ the radius of the droplet, the results are well converged provided a proper scaling for the mobility is used.